Find a number that is 68 greater than three times its opposite
Answer by mananth(16949)

Find a number that is 68 greater than three times its opposite
..
Let the number be x
..
the opposite of this number will be (-x)
..
3 times the opposite of the number = 3*(-x)=-3x
..
add 68
we get
(-3x)+68
x = (-3x+68)
add 3x to both sides
3x+x=68
4x=68
..
divide by 4
4x/4 = 68/4
x= 17 the required number
...
m.ananth@hotmail.ca